# Beware Untrusted Simulators – Reward-Free Backdoor Attacks in Reinforcement Learning

### v2

@@ 1 Introduction
Simulators are a core component of reinforcement learning (RL) and robotics (Makoviychuk et al., 2021; Coumans & Bai, 2016–2021; Todorov et al., 2012), enabling RL approaches for dexterous object manipulation (OpenAI et al., 2018), humanoid robotic control (Xue et al., 2025), and autonomous navigation (Hu et al., 2021). Leveraging simulated interactions, either alone or in combination with real-world data, agents are able to gain task experience in a manner that is both scalable and safe (Da et al., 2025). Therefore, simulators have become a trusted pillar of applied reinforcement learning, with practitioners leveraging external simulators or cloud computing services to train their agents (Bhide et al., 2025).
@@
This trust comes with a cost, however, as it leaves the door open for malicious developers or cloud computing providers to influence the learned behavior of RL policies. Particularly threatening are backdoor attacks in which an agent’s states, transitions, and rewards are manipulated at training time such that they exhibit pre-defined adversarial behavior upon observing a fixed trigger during deployment (Kiourti et al., 2019; Ma et al., 2025; Rathbun et al., 2025; 2024). The backdoor trigger then grants an adversary control over the agent at test time, enabling destructive or biased behavior from an otherwise well-performing agent. Traditional approaches rely on unrealistically invasive assumptions, however, granting the adversary complete control over the training pipeline (Wang et al., 2021) or requiring direct, RAM-level access to the training machine (Rathbun et al., 2024). These assumptions leave a large gap in our understanding of backdoor attacks under more realistic constraints. Therefore, in this work we uncover a previously unexplored threat – malicious developers implanting backdoors into agents by subtly altering the mechanics of their simulator. Such an attack is impossible for traditional backdoor methods to achieve, however, as many simulators, such as the commonly used MuJoCo (Todorov et al., 2012) and Pybullet (Coumans & Bai, 2016–2021) robotics simulators, only receive actions and return environment states, while rewards are computed externally. Therefore this scenario, while realistic, violates a critical assumption of all prior backdoor attacks – the adversary can no longer alter or even observe the agent’s rewards.
@@ 2 Background
In this section we formally define the problem setting, threat model, and adversarial objectives of our attack. We consider two parties: the victim training an RL agent, and the adversary manipulating the agent’s data for malicious purposes. The victim trains an agent to maximize cumulative rewards in a Markov Decision Processes (MDP), typically defined as the tuple $M=(S,A,R,T,\gamma)$ where the state space $S\subset\mathbb{S}$ is a subset of a larger space (e.g. set of all 16 dimensional lidar scans), $A$ is a set of actions, $R:S\times A\rightarrow\mathbb{R}$ is a reward function, $T:S\times A\times S\rightarrow\Delta S$ represents transition probabilities between states given actions, and $\gamma$ discounts the value of rewards over time (Sutton & Barto, 2018). The behavior of an agent is determined by their policy $\pi:S\rightarrow\Delta A$ which represents a probability distribution over actions given states. In this work we consider MDPs with both discrete and continuous action and state spaces $A$. We further consider attacks against Deep Reinforcement Learning (DRL) algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017), Twin Delayed Deep Deterministic policy gradient (TD3) (Fujimoto et al., 2018) and Deep Q-Networks (DQN) (Mnih et al., 2013) as they represent the current state of the art in RL.
@@ 3.1 Reward Free and Universal Backdoor Attacks
where $Q_{\pi}^{M}(s,a)$ represents the expected value of action $a$ in state $s$ given policy $\pi$ and MDP $M$ (Sutton & Barto, 2018). With this we now know that, unless every action has equal value in a given state, there is always some action an agent can choose which performs better than sampling actions at random. We will refer to policies which perform better than random choice as “reasonable” and define them as follows for a given $M$:
@@ 3.3 Daze in Practice
One fortunate consequence of Daze’s design is that it lends itself well to a practical method that is both general and easy to implement. Specifically, since the attack acts in the “inner loop” of the RL training pipeline the attack can be implemented as a wrapper or API around the base MDP $M$ or simulator $\mathcal{S}$. The only requirement for the malicious simulator developer is to choose which trigger and daze functions best suit their test time purposes while minimizing training time detectability. Therefore, in Algorithm 1 we present a practical implementation of Daze as an environment wrapper which is called each time the agent chooses an action $a_{t}$ and looks to sample the next state $s_{t}$.
@@ 4.2 Continuous Action Space Experiments on Real Hardware
In addition to demonstrating the effectiveness of Daze against simulated, continuous action space environments, we further extend our study into policies executing on real, ROS based robots (Quigley, 2009; Diankov & Kuffner, 2008). To this end we design two custom, image-based, continuous control environments using both Turtlebot2 (TUR, 2012) and Fetch (Wise et al., 2016) based agents. Our Turtlebot environment, “Intersection”, is visualized in Figure 1 while our Fetch environment, “Waiter”, is shown in Figure 5. Both “Intersection” and “Waiter” agents, which use depth camera images as observations, are trained in custom Pybullet (Coumans & Bai, 2016–2021) environments designed to be nearly identical to their real world counterparts.
@@
We then transfer our agents, poisoned by Daze, onto real robotic hardware with physical depth camera sensors. Through these results, presented visually on “Intersection” in Figure 1 and “Waiter” in Figure 5, we demonstrate the first example (to the best of our knowledge) of RL backdoor attacks successfully triggering harmful behavior from real, robotic agents. We can also see this concretely in Figure 4, where our “Intersection” agent, poisoned by Daze, exhibits a significant shift in behavior once the trigger is in the environment – following the target action by accelerating forward and almost never turning. This greatly heightens the importance of further research into the prevention and detection of backdoor attacks to mitigate their potentially harmful impacts. In the supplementary material, hosted anonymously here, we provide video demonstrations of our physical experiments. We also include further details for our hardware experiments in the appendix.
@@ 8 Reproducibility Statement
In our appendix and supplementary material (provided anonymously here and uploaded as a zip file to open review) we make sure to provide both an exhaustive list of the hyper-parameters we used along with the code we ran for our experiments. Through this open-sourcing of our experimental setup we look forward to future researchers replicating and building upon our results.
@@ C.4 Hardware Experiment Details
In this section we provide further details on our hardware experiments. Our two real-world scenarios involve a delivery task, “Waiter”, representing robot servers delivering food to guests, and a traffic intersection task, “Intersection”, representing autonomous cars at an intersection. In “Intersection” we used two Kobuki Turtlebot2 robots (TUR, 2012) and in “Waiter” we used a Fetch Robotics Fetch Research Robot (Wise et al., 2016). Simulation and reality are often different with background, lightning, textures, dynamics, etc. This gap introduces many challenges in the sim2real transfer of policies. To alleviate some of the discrepancies, we set up our simulated and real world environments to be as close to each other as possible. Specifically, our simulated and real environments have near identical size and are all surrounded by flat walls to prevent interference from background objects. We can see the similarity of the two environments from the agent’s perspective in Figure 7. Images captured from the agent’s real world sensors were, naturally, more noisy than their simulated counterparts, but we found this had little impact on the performance of the agent’s policy.
@@ E.2 Theorem 1
where we use $V_{*}^{M}(s)$ to denote the optimal value function given state $s$ in MDP $M$ Sutton & Barto (2018). This allows us to compare and transfer policies from $M^{\prime}$ to $M^{\prime}_{b}$, but not necessarily in the opposite direction.

### v3

@@ 1 Introduction
Simulators are a core component of reinforcement learning (RL) and robotics (Makoviychuk et al., 2021; Coumans and Bai, 2016; Todorov et al., 2012), enabling RL approaches for dexterous object manipulation (OpenAI et al., 2018), humanoid robotic control (Xue et al., 2025), and autonomous navigation (Hu et al., 2021). Leveraging simulated interactions, either alone or in combination with real-world data, agents are able to gain task experience in a manner that is both scalable and safe (Da et al., 2025). Therefore, simulators have become a trusted pillar of applied reinforcement learning, with practitioners leveraging external simulators or cloud computing services to train their agents (Bhide et al., 2025).
@@
This trust comes with a cost, however, as it leaves the door open for malicious developers or cloud computing providers to influence the learned behavior of RL policies. Particularly threatening are backdoor attacks in which an agent’s states, transitions, and rewards are manipulated at training time such that they exhibit pre-defined adversarial behavior upon observing a fixed trigger during deployment (Kiourti et al., 2019; Ma et al., 2025; Rathbun et al., 2025; 2024). The backdoor trigger then grants an adversary control over the agent at test time, enabling destructive or biased behavior from an otherwise well-performing agent. Traditional approaches rely on unrealistically invasive assumptions, however, granting the adversary complete control over the training pipeline (Wang et al., 2021) or requiring direct, RAM-level access to the training machine (Rathbun et al., 2024). These assumptions leave a large gap in our understanding of backdoor attacks under more realistic constraints. Therefore, in this work we uncover a previously unexplored threat – malicious developers implanting backdoors into agents by subtly altering the mechanics of their simulator. Such an attack is impossible for traditional backdoor methods to achieve, however, as many simulators, such as the commonly used MuJoCo (Todorov et al., 2012) and Pybullet (Coumans and Bai, 2016) robotics simulators, only receive actions and return environment states, while rewards are computed externally. Therefore this scenario, while realistic, violates a critical assumption of all prior backdoor attacks – the adversary can no longer alter or even observe the agent’s rewards.
@@ 2 Background
In this section we formally define the problem setting, threat model, and adversarial objectives of our attack. We consider two parties: the victim training an RL agent, and the adversary manipulating the agent’s data for malicious purposes. The victim trains an agent to maximize cumulative rewards in a Markov Decision Processes (MDP), typically defined as the tuple $M=(S,A,R,T,\gamma)$ where the state space $S\subset\mathbb{S}$ is a subset of a larger space (e.g. set of all 16 dimensional lidar scans), $A$ is a set of actions, $R:S\times A\rightarrow\mathbb{R}$ is a reward function, $T:S\times A\times S\rightarrow\Delta S$ represents transition probabilities between states given actions, and $\gamma$ discounts the value of rewards over time (Sutton and Barto, 2018). The behavior of an agent is determined by their policy $\pi:S\rightarrow\Delta A$ which represents a probability distribution over actions given states. In this work we consider MDPs with both discrete and continuous action and state spaces $A$. We further consider attacks against Deep Reinforcement Learning (DRL) algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017), Twin Delayed Deep Deterministic policy gradient (TD3) (Fujimoto et al., 2018) and Deep Q-Networks (DQN) (Mnih et al., 2013) as they represent the current state of the art in RL.
@@ 3.1 Reward Free and Universal Backdoor Attacks
where $Q_{\pi}^{M}(s,a)$ represents the expected value of action $a$ in state $s$ given policy $\pi$ and MDP $M$ (Sutton and Barto, 2018). With this we now know that, unless every action has equal value in a given state, there is always some action an agent can choose which performs better than sampling actions at random. We will refer to policies which perform better than random choice as “reasonable” and define them as follows for a given $M$:
@@ 3.3 Daze in Practice
One fortunate consequence of Daze’s design is that it lends itself well to a practical method that is both general and easy to implement. Specifically, the attack can be implemented as API wrapper around the base MDP $M$ or simulator $\mathcal{S}$, however in practice the attack may need to be hidden within normal functionality like domain randomization Tobin et al. (2017). The only requirement for the malicious simulator developer is to choose which trigger and daze functions best suit their test time purposes while minimizing training time detectability. For ease of interpretability, in Algorithm 1, we present a practical implementation of Daze as an environment wrapper which is called each time the agent chooses an action $a_{t}$ and looks to sample the next state $s_{t}$.
@@ 4.2 Continuous Action Space Experiments on Real Hardware
In addition to demonstrating the effectiveness of Daze against simulated, continuous action space environments, we further extend our study into policies executing on real, ROS based robots (Quigley, 2009; Diankov and Kuffner, 2008). To this end we design two custom, image-based, continuous control environments using both Turtlebot2 (1) and Fetch (Wise et al., 2016) based agents. Our Turtlebot environment, “Intersection”, is visualized in Figure 1 while our Fetch environment, “Waiter”, is shown in Figure 5. Both “Intersection” and “Waiter” agents, which use depth camera images as observations, are trained in custom Pybullet (Coumans and Bai, 2016) environments designed to be nearly identical to their real world counterparts.
@@
We then transfer our agents, poisoned by Daze, onto real robotic hardware with physical depth camera sensors. Through these results, presented visually on “Intersection” in Figure 1 and “Waiter” in Figure 5, we demonstrate the first example (to the best of our knowledge) of RL backdoor attacks successfully triggering harmful behavior from real, robotic agents. We can also see this concretely in Figure 4, where our “Intersection” agent, poisoned by Daze, exhibits a significant shift in behavior once the trigger is in the environment – following the target action by accelerating forward and almost never turning. This greatly heightens the importance of further research into the prevention and detection of backdoor attacks to mitigate their potentially harmful impacts. In our github we provide links to video demonstrations of our physical experiments. We also include further details for our hardware experiments in the appendix.
@@ 6 Conclusion
7 Acknowledgments
This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA) under contract W912CG23C0031, by NSF awards CNS-2312875, CNS-2331081, and FMitF 2319500, by an Amazon Research Award and a grant from Coefficient Giving.
@@ 9 Reproducibility Statement
In our appendix and supplementary material (provided on github and uploaded as a zip file to open review) we make sure to provide both an exhaustive list of the hyper-parameters we used along with the code we ran for our experiments. Through this open-sourcing of our experimental setup we look forward to future researchers replicating and building upon our results.
@@ C.4 Hardware Experiment Details
In this section we provide further details on our hardware experiments. Our two real-world scenarios involve a delivery task, “Waiter”, representing robot servers delivering food to guests, and a traffic intersection task, “Intersection”, representing autonomous cars at an intersection. In “Intersection” we used two Kobuki Turtlebot2 robots (1) and in “Waiter” we used a Fetch Robotics Fetch Research Robot (Wise et al., 2016). Simulation and reality are often different with background, lightning, textures, dynamics, etc. This gap introduces many challenges in the sim2real transfer of policies. To alleviate some of the discrepancies, we set up our simulated and real world environments to be as close to each other as possible. Specifically, our simulated and real environments have near identical size and are all surrounded by flat walls to prevent interference from background objects. We can see the similarity of the two environments from the agent’s perspective in Figure 7. Images captured from the agent’s real world sensors were, naturally, more noisy than their simulated counterparts, but we found this had little impact on the performance of the agent’s policy.
@@ E.2 Theorem 1
where we use $V_{*}^{M}(s)$ to denote the optimal value function given state $s$ in MDP $M$ Sutton and Barto (2018). This allows us to compare and transfer policies from $M^{\prime}$ to $M^{\prime}_{b}$, but not necessarily in the opposite direction.